# Proper maps of annuli

### v1

@@ 4. Homogeneous maps via the generic hyperplane rank
Our first result is about rational maps taking two LFT spheres to LFT spheres. The definition of $k_{f}$ for a rational map $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in the projective setting is exactly analogous to the affine setting; it is the smallest dimension $k$ of a $k$-plane in $\mathbb{P}^{N}$ that contains $f(H)$ for a general hyperplane $H\subset\mathbb{P}^{n}$. In the proof we will work in homogeneous coordinates on the projective space, and so a general hyperplane in $H\subset\mathbb{P}^{n}$ is an $n$-dimensional linear subspace in $\mathbb{C}^{n+1}$.
We use $z=(z_{0},\ldots,z_{n})$ for the homogeneous coordinates on $\mathbb{P}^{n}$. To write a rational function $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in homogeneous coordinates we start with $f=\frac{(p_{1},\ldots,p_{N})}{q}$ and order it as $(q,p_{1},\ldots,p_{N})$ and homogenize, that is, we make the denominator the 0th component.
@@
Suppose $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ is a rational map that takes two distinct LFT spheres to two LFT spheres and $k_{f}=N-1$. Then $f\circ\alpha=\beta\circ f$ for some LFTs $\alpha\in\operatorname{Aut}(\mathbb{P}^{n})$ and $\beta\in\operatorname{Aut}(\mathbb{P}^{N})$ neither of which is the identity.
More precisely, using the coordinates above, if one of the spheres on the source (resp. the target) are given by $J_{n}$ (resp. $J_{N}$) and the other by a hermitian matrix $A$ (resp. $B$), then $\alpha$ is represented by the matrix $J_{n}A$ and $\beta$ is represented by the matrix $J_{n}B$.
The main idea behind the proof is that $k_{f}=N-1$ means that the map $f$ induces a map of the corresponding Grassmannians. To figure out the map one uses the Segre variety of the LFT spheres. But now we have two LFT spheres that go to LFT spheres and the second LFT sphere pair means that $f$ induces a different mapping of Grassmannians. Comparing the two maps gives us the corresponding $\alpha$ and $\beta$.
@@
As we said we will work in homogeneous coordinates, that is, in $\mathbb{C}^{n+1}$ and $\mathbb{C}^{N+1}$. Starting with a rational map $f$, we clear denominators and homogenize to represent $f$ by a homogeneous polynomial map $F=(F_{0},F_{1},\ldots,F_{N})\colon\mathbb{C}^{n+1}\to\mathbb{C}^{N+1}$ where $F_{0}$ is the denominator. After applying an LFT we can assume that one of the matrices on both the source and the target is the standard sphere $S^{2n-1}$ and $S^{2N-1}$, that is, it is given by $\mathopen{}\mathclose{{\left\langle J_{n}z,z}}\right\rangle=0$. For any hermitian matrix $M$ and a point $w$, we write the Segre variety of the homogenized LFT sphere given by $\mathopen{}\mathclose{{\left\langle Mz,z}}\right\rangle=0$ as
@@
In affine coordinates, the fact that $f=\frac{p}{q}$ takes the standard sphere to the standard sphere means $\mathopen{}\mathclose{{\left\langle p(z),p(w)}}\right\rangle-q(z)\overline{q(w)}=g(z,\bar{w})\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle z,w}}\right\rangle-1}}\right)$ for a polynomial $g$. In homogeneous coordinates, this translates to
@@
Note that $J_{N}^{-*}=J_{N}$ and $B=B^{*}$, so let $T=J_{N}^{-*}B^{*}=J_{N}B$ and then we have that for all $w$, we have
@@
In particular, $f$ now also takes the LFT sphere $\alpha^{-1}(S^{2n-1})$ to the LFT sphere $\beta^{-1}(S^{2N-1})$. It turns out that these spheres need not be the ones we started with. By iterating, we find that for any $k\in\mathbb{Z}$ we have that
@@
So we also get that $f$ takes $\alpha^{-k}(S^{2n-1})$ to $\beta^{-k}(S^{2N-1})$ for any $k$. In our particular case, when all spheres involved are centered at zero, $f$ takes infinitely many spheres centered at zero of various radii to spheres centered at zero. In particular, in the first step, whenever it takes the standard sphere to the standard sphere and the $r$-sphere to the $R$-sphere, then we find that it also takes the $r^{2}$-sphere to the $R^{2}$-sphere and so on.
@@
Suppose $n\geq 2$ and $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map such that $f(S^{2n-1})\subset S^{2N-1}$ and for some $r,R<1$ we have that $f(rS^{2n-1})\subset RS^{2N-1}$. Suppose also that $k_{f}=N-1$. Then there are two sequences of distinct numbers $\{r_{k}\}$ and $\{R_{k}\}$ converging to zero so that $f(r_{k}S^{2n-1})\subset R_{K}S^{2N-1}$ for every $k$.
@@
We set things up as in the proof of Theorem˜4.1, and so the matrices $A$ and $B$ are given as
@@
We are now ready to prove Theorem˜1.2 using the results of our previous work in [helal-2025-proper].
@@
One direction of the theorem is clear and follows from the discussion so far. What is left to prove is that if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a holomorphic proper map and $k_{f}=N-1$, then $f=UH_{d}$ for some unitary $U$ and an integer $d$. By Theorem˜2.2, $f$ is rational and takes spheres to spheres as in Lemma˜4.2. Therefore, $f$ takes infinitely many spheres centered at zero to spheres centered at zero. In [helal-2025-proper], we have proved that such a rational map must, up to a postcomposition with a unitary, be a direct sum of scaled maps $H_{d}$ for various $d$ or zeros. So for some unitary $U$, we can assume that
(42) $f=U(c_{1}H_{d_{1}}\oplus\cdots c_{\ell}H_{d_{\ell}}\oplus 0),$
where we can assume that $d_{1},\ldots,d_{\ell}$ are distinct. Since we are assuming that $k_{f}=N-1$, we must have that $N_{f}=N$ and hence there are no zero components. Then, via the argument in the proof of Proposition˜2.11, we can see that $k_{f}=N$-$\ell$. Hence $\ell=1$, and $f=UH_{d}$ as required. ∎
Let us now prove Theorem˜1.4 as an application of Theorem˜1.2. We already have that proper holomorphic maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{3,R}$ are rational. If the degree of the map is 1, then via Proposition˜2.7, it is unitarily equivalent to the affine embedding
@@
The theorem will therefore follow if we can show that any map of degree greater than 1 is unitarily equivalent to $H_{2}$. This result follows by Faran’s [faran-1982-maps] classification of proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. The following lemma finishes the proof of Theorem˜1.4.

### v2

@@ 5. Homogeneous maps via the general hyperplane rank
Our first result is about rational maps taking two LFT spheres to LFT spheres. The definition of the general hyperplane rank $k_{f}$ for a rational map $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in the projective setting is exactly analogous to the affine setting; it is the smallest dimension $k$ of a $k$-plane in $\mathbb{P}^{N}$ that contains $f(H)$ for a general hyperplane $H\subset\mathbb{P}^{n}$. In the proof we will work in homogeneous coordinates on the projective space, and so a general hyperplane in $H\subset\mathbb{P}^{n}$ is an $n$-dimensional linear subspace in $\mathbb{C}^{n+1}$.
We use $z=(z_{0},\ldots,z_{n})$ for the homogeneous coordinates on $\mathbb{P}^{n}$. To write a rational map $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in homogeneous coordinates we start with $f=\frac{(p_{1},\ldots,p_{N})}{q}$ and order it as $(q,p_{1},\ldots,p_{N})$ and homogenize, that is, we make the denominator the 0th component.
@@
Suppose $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ is a rational map that takes two distinct LFT spheres to LFT spheres and $k_{f}=N-1$. Then $f\circ\alpha=\beta\circ f$ for some LFTs $\alpha\in\operatorname{Aut}(\mathbb{P}^{n})$ and $\beta\in\operatorname{Aut}(\mathbb{P}^{N})$ neither of which is the identity.
More precisely, using the coordinates above, if one of the spheres on the source (resp. the target) is given by $J_{n}$ (resp. $J_{N}$) and the other by a hermitian matrix $A$ (resp. $B$), then $\alpha$ is represented by the matrix $J_{n}A$ and $\beta$ is represented by the matrix $J_{N}B$.
The main idea behind the proof is that $k_{f}=N-1$ means that the map $f$ induces a map of the corresponding Grassmannians. To figure out the map one uses the Segre variety of the LFT spheres. But now we have two LFT spheres that go to LFT spheres and the second LFT sphere pair means that $f$ induces a different map of Grassmannians. Comparing the two maps gives us the corresponding $\alpha$ and $\beta$.
@@
As we said we will work in homogeneous coordinates, that is, in $\mathbb{C}^{n+1}$ and $\mathbb{C}^{N+1}$. We will write $z$ and $w$ for points in $\mathbb{C}^{n+1}$ and $Z$ and $W$ for points in $\mathbb{C}^{N+1}$. Starting with a rational map $f$, we clear denominators and homogenize to represent $f$ by a homogeneous polynomial map $F=(F_{0},F_{1},\ldots,F_{N})\colon\mathbb{C}^{n+1}\to\mathbb{C}^{N+1}$ where $F_{0}$ is the denominator. After applying an LFT we can assume that one of the matrices on the source corresponds to the standard sphere $S^{2n-1}$, that is, it is given by $\mathopen{}\mathclose{{\left\langle J_{n}z,z}}\right\rangle=0$, and similarly for the target. For any hermitian matrix $M$ and a point $w$, we write the Segre variety of the homogenized LFT sphere given by $\mathopen{}\mathclose{{\left\langle Mz,z}}\right\rangle=0$ as
@@
Write $z=(z_{0},z^{\prime})$ where $z^{\prime}$ refers to the affine coordinates. Then the fact that $f=\frac{p}{q}$ takes the standard sphere to the standard sphere means $\mathopen{}\mathclose{{\left\langle p(z^{\prime}),p(w^{\prime})}}\right\rangle-q(z^{\prime})\overline{q(w^{\prime})}=g(z^{\prime},\bar{w}^{\prime})\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle z^{\prime},w^{\prime}}}\right\rangle-1}}\right)$ for a polynomial $g$. In homogeneous coordinates, this translates to
@@
Note that $J_{N}^{-*}=J_{N}$ and $B=B^{*}$, so let $T=J_{N}^{-*}B^{*}=J_{N}B$ and then we have that for all $w$,
@@
In particular, $f$ now also takes the LFT sphere $\alpha^{-1}(S^{2n-1})$ to the LFT sphere $\beta^{-1}(S^{2N-1})$. It turns out that these spheres need not be the ones we started with. By iterating, we find that for any $k\in\mathbb{Z}$,
@@
So we also get that $f$ takes $\alpha^{-k}(S^{2n-1})$ to $\beta^{-k}(S^{2N-1})$ for any $k$. In our particular case, when all spheres involved are centered at zero, $f$ takes infinitely many spheres centered at zero of various radii to spheres centered at zero. In particular, in the first step, whenever it takes the standard sphere to the standard sphere and the $r$-sphere to the $R$-sphere, we find that it also takes the $r^{2}$-sphere to the $R^{2}$-sphere and so on.
@@
Suppose $n\geq 2$ and $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map such that $f(S^{2n-1})\subset S^{2N-1}$ and for some $r,R<1$ we have that $f(rS^{2n-1})\subset RS^{2N-1}$. Suppose also that $k_{f}=N-1$. Then there are two sequences of distinct numbers $\{r_{k}\}_{k=1}^{\infty}$ and $\{R_{k}\}_{k=1}^{\infty}$ converging to zero so that $f(r_{k}S^{2n-1})\subset R_{k}S^{2N-1}$ for every $k$.
@@
We set things up as in the proof of Theorem˜5.1, and so the matrices $A$ and $B$ are given as
@@
We are now ready to prove Theorem˜1.2 using the results of our previous work in [19].
@@
One direction of the theorem is clear and follows from the discussion so far. What is left to prove is that if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map and $k_{f}=N-1$, then $f=UH_{d}$ for some unitary $U$ and an integer $d$. By Theorem˜2.2, $f$ is rational and takes spheres to spheres as in Lemma˜5.2. Therefore, $f$ takes infinitely many spheres centered at zero to spheres centered at zero. In [19], we have proved that such a rational map must, up to a postcomposition with a unitary, be a direct sum of scaled maps $H_{d}$ for various $d$ or zeros. So for some unitary $U$, we can assume that
(41) $f=U(c_{1}H_{d_{1}}\oplus\cdots\oplus c_{\ell}H_{d_{\ell}}\oplus 0),$
where we can assume that $d_{1},\ldots,d_{\ell}$ are distinct. Since we are assuming that $k_{f}=N-1$, we must have that $N_{f}=N$ and hence there are no zero components. Then, via the argument in the proof of Proposition˜3.4, we can see that $k_{f}=N$-$\ell$. Hence $\ell=1$, and $f=UH_{d}$ as required. ∎
Let us now prove Theorem˜1.4 as an application of Theorem˜1.2. We already have that proper holomorphic maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{3,R}$ are rational. If the degree of the map is 1, then via Proposition˜2.8, it is unitarily equivalent to the affine embedding
@@
The theorem will therefore follow if we can show that any map of degree greater than 1 is unitarily equivalent to $H_{2}$. This result follows by Faran’s [14] classification of proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. The following lemma finishes the proof of Theorem˜1.4.